# A Multi-Simulation Bridge for IoT Digital Twins

### v1

@@ I Introduction
With the widespread adoption of Internet of Things (IoT) and Industrial IoT (IIoT) technologies, Digital Twins (DTs) have evolved into active components—i.e., software entities that maintain continuous bidirectional communication with their physical assets [1, 2, 3]. This paradigm shift allows DTs to provide dynamic digital representations updated through sensor data streams, supporting functionalities such as real-time monitoring, anomaly detection, and adaptive control [4].
In this context, the integration of simulations within their architecture becomes increasingly critical. Simulations would enable DTs to perform “what-if” analyses, optimize performance parameters, and validate control strategies before applying them to the physical asset [5]. The combination of real-time data and virtual experimentation allows DTs to anticipate issues and adapt to changing conditions in complex environments. More in detail, we identified four primary motivations for integrating simulations into DTs:
@@
(1) Enabling predictive and prescriptive capabilities. While IoT-based DTs are effective at capturing the current state of physical systems through real-time data, they are typically limited in projecting future behavior or reasoning under uncertainty. Simulation enhances DTs with predictive capabilities, supporting analysis under varying conditions such as equipment degradation, workload fluctuations, or environmental stressors [6]. This shifts DTs from descriptive digital replicas to prescriptive and prognostic systems capable of anticipating outcomes and triggering proactive interventions.
(2) Coordinating heterogeneous simulations for complex systems. Modeling complex systems often requires the orchestration of multiple, domain-specific simulations, each representing a different aspect of the physical system—e.g., mechanical dynamics, thermal effects, electrical behavior, or environmental interactions [7, 8]. These simulations may operate at different time scales, resolutions, or levels of abstraction and must be integrated to ensure coherence and consistency. This is outside the scope of this paper.
(3) Supporting DT development in the absence of the Physical Twin (PT). In many scenarios – such as early-stage design, safety-critical applications, or high-cost manufacturing – the physical system may not yet exist, be inaccessible, or pose operational risks. In these cases, simulations – e.g., of the IoT system [9] – can act as surrogate PTs, generating mock behaviors and interaction patterns that allow the DT to be developed, tested, and validated in isolation [10].
@@
Despite these promising capabilities, the state of the art reflects only a partial integration of DTs and simulations [11]. This paper presents the DT Simulation Bridge (DT-SB), a software framework designed to integrate DTs with multiple, heterogeneous simulation environments. The DT-SB enables bidirectional data exchange and supports diverse interaction patterns, allowing simulations to dynamically update the DT’s internal state, while DTs can influence and tune simulation parameters in real time. Unlike traditional co-simulation frameworks [12], the DT-SB connects a DT with multiple, independent simulations and acts solely as an intermediary. The coordination and aggregation of simulation results are delegated entirely to the DT, preserving architectural decoupling and maximizing flexibility. We describe the modular architecture and core software components of the DT-SB, emphasizing its interoperability with diverse simulation models and protocols. An experimental evaluation demonstrates the bridge’s effectiveness in accelerating iterative design cycles, facilitating virtual commissioning, and enabling live behavioral analysis under realistic operational conditions.
@@ II Related Work
Simulation is a key technology in DTs, envisioned as a tool that can be adopted across all phases of the PT lifecycle [5]. Despite the strong overlap between simulation and DTs making it difficult to draw a line between the technologies [11], it is becoming more and more clear that, especially in the IoT domain, DTs extend the capability of “static” simulations by incorporating real-time data [13].
@@
The DT Simulation Bridge integrates and orchestrates multiple simulations with different models in a coherent framework. At a surface level, it is then similar to distributed simulation [14] and co-simulation [12]. There are, though, fundamental differences in the approach. Distributed simulation usually concerns the execution of a single complex simulation on distributed nodes. Differently, co-simulation explores the possibility of composing different simulation models of individual parts to simulate a complex system and coordinate the progress of time between the different independent simulation units. The purpose of the DT Simulation Bridge is, instead, to connect one DT with multiple simulations. The coordination and combination of such simulation results are expected to be managed by the DT and not by the bridge itself, which only acts as an intermediary. Nevertheless, despite these differences in purposes, there are some significant aspects from these two related areas which influence the design of the bridge.
Integrating simulations requires standardising the interaction across components. For distributed simulation, the High-Level Architecture (HLA) defines a set of constraints to design interoperable federations of simulations which can be executed on a distributed runtime infrastructure [15]. Despite in our case simulations do not need to exchange data between each other, the need for a standardized interface is essential to allow the DT to seamlessly interact with multiple different simulations. Similarly, in the co-simulation domain, the Functional Mockup Interface (FMI) is an industrial de facto standard [16] to bridge different simulation models. There, models of different nature may declare connection points to create input/output chains which combine the output of individual solvers. We follow this principle in the design of the DT-SB to support simulation models which can be updated over time with the data that is captured by the DT itself.
@@ III Simulation Bridge Interaction Patterns
Figure 2: Architecture of the Simulation Bridge: (a) modular components supporting DT interaction and developer-driven PT simulation, and (b) communication flow between DTs, the bridge, and simulators for simulation control and integration.
@@ V-A Matlab Agent
The MATLAB Agent ($SA_{\text{MATLAB}}$) represents a decoupled simulation agent that integrates MATLAB’s computational capabilities within the DT-SB architecture. The agent communicates with the DT-SB using the AMQP protocol, ensuring reliable and platform-agnostic messaging. Four distinct integration strategies were investigated, each offering different trade-offs in terms of latency, flexibility, and architectural coupling.
1. File-Based Exchange. In this approach, the simulator receives simulation inputs via files, executes the model, and stores the results in an output file. The $SA_{\text{MATLAB}}$ reads the output file, encapsulates its contents as the response payload, and forwards the message to the DT-SB. This method offers high architectural decoupling but incurs substantial input/output overhead, making it unsuitable for streaming simulations due to poor performance and high latency.
2. TCP/IP Communication. This approach employs network-based IPC, enabling real-time interaction between the simulator and the $SA_{\text{MATLAB}}$. The simulator sends result data directly to the $SA_{\text{MATLAB}}$ over a TCP/IP socket, formatted as JSON. Each result corresponds to a single message in the stream, making this method highly compatible with streaming simulation modes. The use of network IPC ensures that the MA and simulator remain decoupled at the operating system process level, supporting distributed deployment.
3. Shared Memory IPC. In this configuration, the $SA_{\text{MATLAB}}$ and MATLAB exchange data via shared memory segments. Simulation requests and responses are written and read directly from volatile memory, resulting in extremely low latency and high throughput. However, this approach requires explicit synchronization using mutexes or locks to avoid race conditions and deadlocks. Additionally, shared memory access necessitates strict control over memory layout, reducing flexibility.
4. MATLAB Engine API. This method utilizes the MATLAB Engine API to directly invoke MATLAB functions from Python. While it simplifies invocation of MATLAB computations, it imposes structural constraints on model definitions and requires the simulator to conform to specific function-call semantics, thus reducing modularity and reusability.
A comprehensive evaluation of these four approaches has been conducted to determine the most suitable integration strategy for batch and streaming simulation modes. Although shared memory IPC achieves the lowest latency, it required rigid memory management and offers limited flexibility. Ultimately, the Matlab Engine API approach has been selected for batch simulation. The TCP/IP approach has been selected as the preferred integration method for streaming mode simulation due to its balance of modularity, ease of integration, and suitability for distributed architectures.
@@ VI Experimental Evaluation
To evaluate the performance of the proposed DT-SB and its integration with simulation agents, we conducted a series of experiments focusing on protocol overhead and agent-level performance. All tests were executed on an Apple MacBook Air (2023) equipped with an Apple Silicon M2 processor and 8 GB of unified memory. MATLAB was natively installed on macOS, and Python profiling modules were used to gather performance metrics.
@@ VI-A Simulation Bridge Overhead
Figure 3: AnyLogic PT simulation.
VI-B Physical Twin Simulation
The simulated environment was developed using the DT-SB integrated agent for AnyLogic to support DT creation for a microfactory scenario, featuring an indexed line with a conveyor, workstations, and a robot handling parts between loading areas, as shown in Figure 3.
The two-dimensional model, includes conveyors reporting status, photocells detecting product passage, and transfer stations with custom switches emulating pushers. Workstations have virtual actuators reporting status, and the central robot, modeled via the Robot element, manages item positioning and state updates. The simulation integrates all key entities and focuses on assessing the communication overhead of bidirectional data exchange with the DT via the DT-SB.
Process logic uses a block diagram: agents are generated by a Source block, controlled by a Hold block linked to UI buttons, processed by a ProcessByRobot block, moved via a Convey block, and removed by a Sink block. The DT processes data as if from real PTs. The measured communication overhead from AnyLogic to the DT via MQTT averaged $6.5\,\mathrm{ms}$ with a standard deviation of $10\,\mathrm{ms}$.
This setup offers a controlled, repeatable environment for testing DTs, supporting experimentation with PT conditions, communication protocols, synchronization, and CPS integration strategies.
@@ VII Conclusion
In this paper, we tackled the integration of heterogeneous simulations with DTs via a general-purpose middleware framework. We analyzed key interaction patterns—batch, streaming, and PT simulation—supporting various DT lifecycle use cases. We proposed the DT Simulation Bridge (DT-SB) architecture and a lightweight protocol that ensures a uniform interface between DTs and diverse simulation environments. The DT-SB enables data translation, protocol adaptation, and interoperability across platforms. We validated our solution through two case studies: (i) MATLAB-based streaming performance evaluation; and (ii) AnyLogic-based PT simulation of a cyber-physical scenario. Future work will focus on classifying integration patterns, extending to co-simulation and hybrid setups, refining synchronization, and exploring adaptive orchestration in distributed environments.

### v2

@@ I Introduction
With the widespread adoption of Internet of Things (IoT) and Industrial IoT (IIoT) technologies, Digital Twins (DTs) have evolved into active components—i.e., software entities that maintain continuous bidirectional communication with their physical assets [grieves2014digital, dt_manufacturing_review, dt_industry_statoftheart]. This paradigm shift allows DTs to provide dynamic digital representations updated through sensor data streams, supporting real-time monitoring, anomaly detection, and adaptive control [Cimino_2019].
In this context, integrating simulations into the DT architecture becomes increasingly critical. Simulations would enable DTs to perform “what-if” analyses, optimize performance parameters, and validate control strategies before applying them to the physical asset [Boschert2016]. The combination of real-time data and virtual experimentation allows DTs to anticipate issues and adapt to changing conditions in complex environments. More in detail, we identified four primary motivations for integrating simulations into DTs:
@@
(1) Enabling predictive and prescriptive capabilities. While IoT-based DTs effectively capture the current state of physical systems through real-time data, open challenges remain in enhancing their ability to forecast future behavior and reason under uncertainty [10015420]. Simulation enhances DTs with predictive capabilities, supporting analysis under varying conditions such as equipment degradation, workload fluctuations, or environmental stressors [KLEIN2026103063]. This shifts DTs from descriptive digital replicas to prescriptive and prognostic systems capable of anticipating outcomes and triggering proactive interventions.
(2) Coordinating heterogeneous simulations for complex systems. Modeling complex systems often requires the orchestration of multiple, domain-specific simulations, each representing a different aspect of the physical system—e.g., mechanical dynamics, thermal effects, electrical behavior, or environmental interactions [10.1115/1.4049861, Fitzgerald2019]. These simulations may operate at different time scales, resolutions, or levels of abstraction and must be integrated to ensure coherence and consistency. This is outside the scope of this paper.
(3) Supporting DT development in the absence of the Physical Twin (PT). In many scenarios – such as early-stage design, safety-critical applications, or high-cost manufacturing – the physical system may not yet exist, be inaccessible, or pose operational risks. In these cases, simulations – e.g., of the IoT system [7568309] – can act as surrogate PTs, generating mock behaviors and interaction patterns that allow the DT to be developed, tested, and validated in isolation [AITALLA20191331, 11087838].
@@
Despite these promising capabilities, the state of the art reflects only a partial integration of DTs and simulations [WOOLEY2023940]. This paper presents the DT Simulation Bridge (DT-SB), a software framework designed to integrate DTs with multiple, heterogeneous simulation environments. The DT-SB enables bidirectional data exchange and supports diverse interaction patterns, allowing simulations to dynamically update the DT’s internal state, while DTs can influence and tune simulation parameters in real time. Unlike traditional co-simulation frameworks [Gomes&18], the DT-SB connects a DT with multiple, independent simulations and acts solely as an intermediary. The coordination and aggregation of simulation results are delegated entirely to the DT, preserving architectural decoupling and maximizing flexibility. We describe the modular architecture and core software components of the DT-SB, emphasizing its interoperability with diverse simulation models and protocols. An experimental evaluation demonstrates the bridge’s effectiveness in accelerating iterative design cycles, facilitating virtual commissioning, and enabling live behavioral analysis under realistic operational conditions.
@@ II Related Work
Simulation is a key technology in DTs, envisioned as a tool that can be adopted across all phases of the PT lifecycle [Boschert2016]. Despite the strong overlap between simulation and DTs making it difficult to draw a line between the technologies [WOOLEY2023940], it is becoming more and more clear that, especially in the IoT domain, DTs extend the capability of “static” simulations by incorporating real-time data [PHANDEN2021174].
@@
The DT Simulation Bridge integrates and orchestrates multiple simulations with different models in a coherent framework. At a surface level, it is then similar to distributed simulation [2000daglib] and co-simulation [Gomes&18]. There are, though, fundamental differences in the approach. Distributed simulation usually concerns the execution of a single complex simulation on distributed nodes. Differently, co-simulation explores the possibility of composing different simulation models of individual parts to simulate a complex system and coordinate the progress of time between the different independent simulation units. The purpose of the DT Simulation Bridge is, instead, to connect one DT with multiple simulations. The coordination and combination of such simulation results are expected to be managed by the DT and not by the bridge itself, which only acts as an intermediary. Nevertheless, despite these differences in purposes, there are some significant aspects from these two related areas which influence the design of the bridge.
Integrating simulations requires standardising the interaction across components. For distributed simulation, the High-Level Architecture (HLA) defines a set of constraints to design interoperable federations of simulations which can be executed on a distributed runtime infrastructure [568652]. Despite in our case simulations do not need to exchange data between each other, the need for a standardized interface is essential to allow the DT to seamlessly interact with multiple different simulations. Similarly, in the co-simulation domain, the Functional Mockup Interface (FMI) is an industrial de facto standard [FMIStandard2.0] to bridge different simulation models. There, models of different nature may declare connection points to create input/output chains which combine the output of individual solvers. We follow this principle in the design of the DT-SB to support simulation models which can be updated over time with the data that is captured by the DT itself.
@@ III Simulation Bridge Interaction Patterns
Figure 2: Architecture of the Simulation Bridge with its main components and the associated interaction flows.
@@ V Simulation Agents
This section presents the simulation agents integrated into the DT-SB architecture, focusing on the MATLAB and the AnyLogic Agents enabling the execution of computational and simulation models while maintaining modularity and seamless communication with the DT-SB.
@@ V-A Matlab Agent
The MATLAB Agent ($SA_{\text{MATLAB}}$) serves as a decoupled simulation component integrating MATLAB’s computational capabilities within the DT-SB architecture. It communicates with the DT-SB through the AMQP protocol, enabling reliable and platform-independent message exchange. Four integration strategies were explored, each balancing latency, flexibility, and architectural coupling differently.
In the file-based approach, simulation inputs are provided as files, and results are written back to disk for the agent to read and forward to the DT-SB. This method ensures strong decoupling but introduces considerable input/output overhead, making it unsuitable for real-time or streaming simulations. The TCP/IP communication approach enables real-time data exchange via socket-based interprocess communication, where results are transmitted as JSON messages. This solution supports continuous streaming and distributed deployment with moderate latency. A third strategy relies on shared memory interprocess communication, where requests and responses are exchanged directly through memory segments. While this configuration achieves minimal latency and high throughput, it requires explicit synchronization and strict control over memory layout, reducing flexibility. Finally, the MATLAB Engine API allows direct invocation of MATLAB functions from Python, simplifying model execution but imposing tighter structural constraints and lowering modularity.
Experimental evaluation showed that shared memory communication provides the lowest latency but lacks adaptability. Therefore, the MATLAB Engine API was selected for batch simulations, whereas the TCP/IP approach was adopted for streaming mode, offering a balanced trade-off between performance, modularity, and ease of integration.
@@ V-B AnyLogic Agent
(a) The UR10e robotic arm in MatLab.
(b) AnyLogic Microfactory simulation.
Figure 3: Simulated Physical Twins integrated with the Simulation Bridge.
@@ VI Experimental Evaluation
To evaluate the performance of the proposed DT-SB and its integration with simulation agents, we conducted a series of experiments focusing on protocol overhead and agent-level performance on commercial computing hardware (MacBook Air Apple Silicon M2 processor and 8 GB of unified memory) with Python implementation of the proposed DT-SB  [^3] Simulation Bridge GitHub Repo: https://github.com/INTO-CPS-Association/simulation-bridge.
@@ VI-A Simulation Bridge Overhead
Anylogic Agent Overhead. To evaluate the performance of the AnyLogic agent a dedicated testing campaign was conducted to measure the system’s intrinsic overhead, independently of simulation model complexity aligned with the MATLAB agent analysis. The simulation used for the experiment is shown in Figure 3b. The AnyLogic model employed is event-based, meaning that the simulation flow is internally controlled and does not rely on a fixed time step, unlike Table III, which reports the MATLAB agent’s overhead at different frequencies. To ensure comparability with previous analyses, the simulation was executed at various run-time speeds (1x, 2x, 5x, 10x, 25x, 50x). The results show an extremely low and stable average overhead, ranging from 3.4 ms to 8.6 ms across different streaming configurations, as reported in Table IV. These findings confirm that $SA_{\text{AnyLogic}}$ introduces negligible latency, making it suitable for real-time streaming applications.
TABLE IV: AnyLogic Agent’s overhead at different execution speeds.
Execution Speed 1x 2x 5x 10x 25x 50x
Agent Overhead (ms) 8.6 4.9 8.3 3.4 5.3 3.6
VI-B Physical Twin Simulations
To validate and analyze the proposed Simulation Bridge framework and the integration with DTs, two simulation environments were developed to emulate distinct aspects of the PTs behaviors: one in MATLAB/Simulink, focusing on robotic motion control and kinematics, and one in AnyLogic, modeling a microfactory scenario for system-level DT integration and communication assessment.
The first simulation involves a Universal Robots UR10e robotic arm executing a smooth point-to-point motion in joint space as depicted in Figure 3a. The objective is to move the robot from an initial joint configuration to a target configuration within a specified time, following an optimally smooth trajectory and employing a simple closed-loop controller to track this trajectory. To ensure smooth motion, a fifth-order (quintic) polynomial trajectory was defined for each of the robot’s six joints. A quintic polynomial has six coefficients, allowing the imposition of six boundary conditions (initial and final position, velocity, and acceleration). This ensures continuity of the first and second derivatives—velocity and acceleration—throughout the entire motion.
The robotic simulation was implemented using the Robotics System Toolbox and a rigid-body model of the UR10e with gravity enabled. At each time step, the control law computed a new joint velocity command, and joint angles were updated via simple Euler integration. This iterative process continued until the planned trajectory duration elapsed. Completion was detected when the final joint configuration error dropped below a predefined negligible threshold, ensuring accurate convergence to the target configuration.
The second simulated environment was developed using the DT-SB integrated agent within AnyLogic to support DT creation for a Microfactory scenario. The model represents an indexed production line comprising a conveyor, multiple workstations, and a robotic arm responsible for handling parts between loading and unloading areas, as illustrated in Figure 3b.
The two-dimensional model includes conveyors capable of reporting operational status, photocells detecting product passage, and transfer stations with custom switches emulating pusher mechanisms. Workstations are equipped with virtual actuators that communicate their operational state, while the central robot, modeled via the Robot element, manages item positioning and state updates. Process logic is implemented using a block-diagram approach: agents are generated by a Source block, controlled by a Hold block linked to user interface commands, processed by a ProcessByRobot block, transferred using a Convey block, and finally removed by a Sink block.
Together, these simulation environments provided a controlled and complementary validation setup for the proposed approach. The MATLAB/Simulink model enabled low-level evaluation of motion control and trajectory tracking, while the AnyLogic microfactory scenario assessed system-level communication, synchronization, and interoperability. This combination demonstrated the framework’s flexibility in connecting diverse simulation domains within a unified DT integration architecture.
@@ VII Conclusion
We propose the DT Simulation Bridge (DT-SB), a lightweight middleware architecture and protocol that unifies data exchange and interoperability between digital twins and heterogeneous simulation environments. The DT-SB enables seamless integration across batch, streaming, and PT simulation interaction patterns, supporting diverse DT lifecycle applications. The DT-SB handles data translation, protocol adaptation, and synchronization between DTs and external simulators. Its effectiveness was demonstrated through two case studies based on MATLAB and AnyLogic simulations of cyber-physical scenarios. Future work will extend this approach toward co-simulation and hybrid setups, refined synchronization mechanisms, and adaptive orchestration in distributed environments together with the integration with DT open-source frameworks [PICONE2021100661].